Use double integration to find the volume of the solid bounded by the paraboloid z= 9x^2 + y^2, below by the plane z = 0, and laterally by the planes x = 0, y = 0, x = 3, and y = 2.


1
Expert's answer
2021-10-21T14:46:20-0400
"V=\displaystyle\int_{0}^3\displaystyle\int_{0}^2(9x^2+y^2)dydx"

"=\displaystyle\int_{0}^3[9x^2y+y^3/3]\begin{matrix}
 2\\
 0
\end{matrix}dx""=\displaystyle\int_{0}^3(18x^2+8/3)dx"

"=[6x^3+8x/3]\begin{matrix}
 3 \\
 0
\end{matrix}"

"=170 (cubic\ units)"